Find h'(1) given that f(1) = -5, f '(1) = 2, g(1) = 3, and g'(1) = 2

h(x) = f(x)/g(x)

h'(1) =
1
Expert's answer
2011-03-14T08:28:11-0400
If h(x) = f(x)/g(x), then
h'(x) = {f'(x)g(x) - g'(x)f(x)}/g2(x)
Therefore
h'(1) = {f'(1)g(1) - g'(1)f(1)}/g2(1) = {2*3 - 2*(-5)}/32 = 16/9